# Not All Data are Good Labels: On the Self-supervised Labeling for Time Series Forecasting

### v3

@@ Appendix F Related Work Revisit
Channel-independent Transformers. Time Series Transformers (TST) [53, 40, 54] have recently led to significant progress in the TSF problem, demonstrating convincing superiority over traditional methods and convolution-based models. Inspired by [48], [27] incorporates the Channel Independent (CI) design (sharing weights across all channels) to introduce PatchTST, a new state-of-the-art (SOTA) model that significantly benefits from CI and the patching operation. Multiple works follow this practice and achieve excellent performance in TSF [6, 19, 50].
Channel-wise Transformers. Building on TST, more recent research has focused on designing Transformers capable of capturing channel dependencies inherent in multivariate time series data. Notable examples include Crossformer [51], iTransformer (iTrans[21], Dsformer [47], and card [37]. These models shed light on exploiting inter-series relationships to improve forecasting accuracy. Furthermore, having observed that Time Series Transformers are inherently unstable due to a sharp loss scape, [10] propose a sharpness-aware optimizer to mitigate such issues. Their work focuses on an optimization-level approach involving a two-step backward. Nonetheless, channel-wise Transformers still suffer from overfitting on small datasets. By sharpness analysis, our proposed Scam locates the overfitting issue and provides a solution from a data-level perspective.
Linear or MLP-based Models. In contrast to the quadratic complexity of Transformers, lightweight linear or MLP-based models have emerged as competitive alternatives offering simplicity and efficiency. RLinear and RMlp [15] verify that a vanilla linear model or a 2-layer MLP, when combined with a widely-adopted normalization method [12], can achieve near SOTA performance in TSF. Further research [52] on channel dependencies within linear and MLP-based models has yielded performance improvements over previous CI approaches. Moreover, the models [44, 45] that directly learn linear regression or MLP-based models on complex frequency features have achieved remarkable performances. Most recently, SparseTSF [16], a highly lightweight model, incorporates 1D convolutions as down-sampling modules and learns linear parameters on the down-sampled values of the original series. cycleNet [17], a SOTA model that explicitly captures periodic trend features to enhance vanilla linear or MLP-based models to be on par with Transformer-based models.
@@ G.1 Datasets
  • ETT [53]: This dataset contains 7 factors of electricity transformers, recorded between July 2016 and July 2018. The subsets ETTh1 and ETTh2 are recorded hourly, while ETTm1 and ETTm2 are recorded every 15 minutes.
  • Electricity [40]: This dataset records the hourly electricity consumption of 321 clients.
  • Traffic [40]: This dataset collects hourly road occupancy rates measured by 862 sensors across the San Francisco Bay Area freeways, spanning from January 2015 to December 2016.
  • Weather [40]: This dataset includes 21 meteorological factors, recorded every 10 minutes at the Weather Station of the Max Planck Biogeochemistry Institute in 2020.
  • PeMS: This dataset contains public traffic network data from California, collected at 5-minute intervals. We use the same four subsets (PeMS03, PeMS04, PeMS07, PeMS08) as adopted in iTransformer (iTrans[21].
For the ETT datasets, we divide them by ratio $\{0.6,0.2,0.2\}$ into train set, validation set, and test set. For Electricity, Traffic, and Weather, we follow the same split ratio of $\{0.7,0.1,0.2\}$ as in TimesNet [41, 53, 40]. For the PeMS datasets, we split them using the ratio $\{0.6,0.2,0.2\}$ following the same setting as iTrans [21]. All datasets are scaled using the mean and variance of their respective training sets, a standard practice in TSF [41]. The statistics of all used datasets are listed in Table 4.
@@ G.2 Backbones
    Mlp [15], a 2-layer MLP model combined with Reversible Instance Normalization [12].
@@
    cycleNet [17], a 2-layer MLP equipped with efficient cycle modeling that belongs to a general seasonal-trend decomposition method (Cycle/MLP in the original paper).
@@
    iTrans [21], a Transformer-based model that computes attention scores on the inverted series along the channel dimension.
@@
    PatchTST [27], a Channel-Independent Transformer-based model that uses patching to tokenize the input.
@@ G.3 Reproductivity
Hyperparameters and Settings. All experiments and methods are implemented in Python and PyTorch [29] and conducted on two Nvidia RTX A5000 Ada generation GPUs (32GB VRAMs) and two Nvidia RTX A6000 GPUs (48GB VRAMs). We use the ADAM optimizer [13] with a learning rate initialized as $\eta=0.001$ for all settings. Unlike prior works [21, 16] that set a fixed, small number of training epochs, we adopt an early stopping strategy based on the MSE metric of the validation set, with a patience of 20 epochs.
@@
Rerun Baselines as Additional Comparisons. To ensure fair comparisons with baseline results, we include both the results reported in the original papers and the results from our own re-implementations. Specifically: The Mlp and cycleNet results are taken from cycleNet paper [17]. The PatchTST and iTrans results are from the iTrans paper [21], as the PatchTST paper [27] does not provide results with a look-back length of 96.
@@
The Result Discrepancy on PeMS Datasets. We also evaluate the larger traffic dataset, PeMS, which has been previously examined in both iTrans and cycleNet. However, we observe major discrepancies between our results and those reported in the original papers [21, 17].
@@ Appendix H More Experiment Results
Table 8: Full results for performance comparisons between backbone models and their integration with our proposals on ETT, Electricity, Traffic, and Weather datasets. The results for Mlp and cycleNet are taken from the cycleNet paper [27], while the results for PatchTST and iTrans are from the iTransformer paper [21].

### v4

@@ Appendix F Related Work Revisit
Channel-independent Transformers. Time Series Transformers (TST) [informer, autoformer, fedformer] have recently led to significant progress in the TSF problem, demonstrating convincing superiority over traditional methods and convolution-based models. Inspired by [dlinear], [patchtst] incorporates the Channel Independent (CI) design (sharing weights across all channels) to introduce PatchTST, a new state-of-the-art (SOTA) model that significantly benefits from CI and the patching operation. Multiple works follow this practice and achieve excellent performance in TSF [timesfm, unitime, mtst].
Channel-wise Transformers. Building on TST, more recent research has focused on designing Transformers capable of capturing channel dependencies inherent in multivariate time series data. Notable examples include Crossformer [crossformer], iTransformer (iTrans[itrans], Dsformer [dsformer], and card [card]. These models shed light on exploiting inter-series relationships to improve forecasting accuracy. Furthermore, having observed that Time Series Transformers are inherently unstable due to a sharp loss scape, [samformer] propose a sharpness-aware optimizer to mitigate such issues. Their work focuses on an optimization-level approach involving a two-step backward. Nonetheless, channel-wise Transformers still suffer from overfitting on small datasets. By sharpness analysis, our proposed Scam locates the overfitting issue and provides a solution from a data-level perspective.
Linear or MLP-based Models. In contrast to the quadratic complexity of Transformers, lightweight linear or MLP-based models have emerged as competitive alternatives offering simplicity and efficiency. RLinear and RMlp [rlinear] verify that a vanilla linear model or a 2-layer MLP, when combined with a widely-adopted normalization method [revin], can achieve near SOTA performance in TSF. Further research [lift] on channel dependencies within linear and MLP-based models has yielded performance improvements over previous CI approaches. Moreover, the models [fits, frequency] that directly learn linear regression or MLP-based models on complex frequency features have achieved remarkable performances. Most recently, SparseTSF [sparsetsf], a highly lightweight model, incorporates 1D convolutions as down-sampling modules and learns linear parameters on the down-sampled values of the original series. cycleNet [cyclenet], a SOTA model that explicitly captures periodic trend features to enhance vanilla linear or MLP-based models to be on par with Transformer-based models.
@@ G.1 Datasets
  • ETT [informer]: This dataset contains 7 factors of electricity transformers, recorded between July 2016 and July 2018. The subsets ETTh1 and ETTh2 are recorded hourly, while ETTm1 and ETTm2 are recorded every 15 minutes.
  • Electricity [autoformer]: This dataset records the hourly electricity consumption of 321 clients.
  • Traffic [autoformer]: This dataset collects hourly road occupancy rates measured by 862 sensors across the San Francisco Bay Area freeways, spanning from January 2015 to December 2016.
  • Weather [autoformer]: This dataset includes 21 meteorological factors, recorded every 10 minutes at the Weather Station of the Max Planck Biogeochemistry Institute in 2020.
  • PeMS: This dataset contains public traffic network data from California, collected at 5-minute intervals. We use the same four subsets (PeMS03, PeMS04, PeMS07, PeMS08) as adopted in iTransformer (iTrans[itrans].
For the ETT datasets, we divide them by ratio $\{0.6,0.2,0.2\}$ into train set, validation set, and test set. For Electricity, Traffic, and Weather, we follow the same split ratio of $\{0.7,0.1,0.2\}$ as in TimesNet [timesnet, informer, autoformer]. For the PeMS datasets, we split them using the ratio $\{0.6,0.2,0.2\}$ following the same setting as iTrans [itrans]. All datasets are scaled using the mean and variance of their respective training sets, a standard practice in TSF [timesnet]. The statistics of all used datasets are listed in Table 4.
@@ G.2 Backbones
    Mlp [rlinear], a 2-layer MLP model combined with Reversible Instance Normalization [revin].
@@
    cycleNet [cyclenet], a 2-layer MLP equipped with efficient cycle modeling that belongs to a general seasonal-trend decomposition method (Cycle/MLP in the original paper).
@@
    iTrans [itrans], a Transformer-based model that computes attention scores on the inverted series along the channel dimension.
@@
    PatchTST [patchtst], a Channel-Independent Transformer-based model that uses patching to tokenize the input.
@@ G.3 Reproductivity
Hyperparameters and Settings. All experiments and methods are implemented in Python and PyTorch [pytorch] and conducted on two Nvidia RTX A5000 Ada generation GPUs (32GB VRAMs) and two Nvidia RTX A6000 GPUs (48GB VRAMs). We use the ADAM optimizer [adam] with a learning rate initialized as $\eta=0.001$ for all settings. Unlike prior works [itrans, sparsetsf] that set a fixed, small number of training epochs, we adopt an early stopping strategy based on the MSE metric of the validation set, with a patience of 20 epochs.
@@
Rerun Baselines as Additional Comparisons. To ensure fair comparisons with baseline results, we include both the results reported in the original papers and the results from our own re-implementations. Specifically: The Mlp and cycleNet results are taken from cycleNet paper [cyclenet]. The PatchTST and iTrans results are from the iTrans paper [itrans], as the PatchTST paper [patchtst] does not provide results with a look-back length of 96.
@@
The Result Discrepancy on PeMS Datasets. We also evaluate the larger traffic dataset, PeMS, which has been previously examined in both iTrans and cycleNet. However, we observe major discrepancies between our results and those reported in the original papers [itrans, cyclenet].
@@ Appendix H More Experiment Results
Table 8: Full results for performance comparisons between backbone models and their integration with our proposals on ETT, Electricity, Traffic, and Weather datasets. The results for Mlp and cycleNet are taken from the cycleNet paper [patchtst], while the results for PatchTST and iTrans are from the iTransformer paper [itrans].